# pySLAM: An Open-Source, Modular, and Extensible Framework for SLAM

### v2

@@ Introduction
When you run the script main_slam.py (main_map_dense_reconstruction.py): - You can save the current map state by pressing the button Save on the GUI. This saves the current map along with front-end, and backend configurations into the default folder results/slam_state (results/slam_state_dense_reconstruction). - To change the default saving path, open config.yaml and update target folder_path in the section: bash   SYSTEM_STATE:     folder_path: results/slam_state   # default folder path (relative to repository root) where the system state is saved or reloaded
@@
    . pyenv-activate.sh   #  Activate pyslam python virtual environment. This is only needed once in a new terminal.
    ./main_map_viewer.py  #  Use the --path options to change the input path
@@
    SYSTEM_STATE:
      load_state: True                  # flag to enable SLAM state reloading (map state + loop closing state)
      folder_path: results/slam_state   # default folder path (relative to repository root) where the system state is saved or reloaded
Note that pressing the Save button saves the current map, front-end, and backend configurations. Reloading a saved map overwrites the current system configurations to ensure descriptor compatibility.
@@
Estimated trajectories can be saved in three different formats: TUM (The Open Mapping format), KITTI (KITTI Odometry format), and EuRoC (EuRoC MAV format). pySLAM saves two types of trajectory estimates:
  • Online: In online trajectories, each pose estimate depends only on past poses. A pose estimate is saved at the end of each front-end iteration on current frame.
  • Final: In final trajectories, each pose estimate depends on both past and future poses. A pose estimate is refined multiple times by LBA windows that cover it and by GBA during loop closures.
@@
SAVE_TRAJECTORY:
  save_trajectory: True
  format_type: kitti             # supported formats: `tum`, `kitti`, `euroc`
  output_folder: results/metrics # relative to pyslam root folder
  basename: trajectory           # basename of the trajectory saving output
SLAM GUI
@@
  • Save: Save the map into the file map.json. You can visualize it back by using the script /main_map_viewer.py (as explained above).
@@
  • Draw Ground Truth: If a ground truth dataset is loaded (e.g., from KITTI, TUM, EUROC, or REPLICA), you can visualize it by pressing this button. The ground truth trajectory will be displayed in 3D and progressively aligned (approximately every 30 frames) with the estimated trajectory. The alignment improves as more samples are added to the estimated trajectory. After ~20 frames, if the button is pressed, a window will appear showing the Cartesian alignment errors (ground truth vs. estimated trajectory) along the axes.
Monitor the logs for tracking, local mapping, and loop closing simultaneously
The logs generated by the modules local_mapping.py, loop_closing.py, loop_detecting_process.py, and global_bundle_adjustments.py are collected in the files local_mapping.log, loop_closing.log, loop_detecting.log, and gba.log, which are all stored in the folder logs. For debugging, you can monitor a parallel flow by running the following command in a separate shell:
@@
Press CTRL+A and then CTRL+Q to exit from tmux environment.
Supported components and models
Supported local features
At present time, the following feature detectors are supported:
@@
For more information, refer to local_features/feature_types.py file. Some of the local features consist of a joint detector-descriptor. You can start playing with the supported local features by taking a look at test/cv/test_feature_manager.py and main_feature_matching.py.
In both the scripts main_vo.py and main_slam.py, you can create your preferred detector-descritor configuration and feed it to the function feature_tracker_factory(). Some ready-to-use configurations are already available in the file local_features/feature_tracker.configs.py
The function feature_tracker_factory() can be found in the file local_features/feature_tracker.py. Take a look at the file local_features/feature_manager.py for further details.
@@
Supported matchers
See the file local_features/feature_matcher.py for further details.
Supported global descriptors and local descriptor aggregation methods
@@
Different loop closing methods are available. These combines the above aggregation methods and global descriptors. See the file loop_closing/loop_detector_configs.py for further details.
Supported depth prediction models
@@
Supported volumetric mapping methods
@@
  • TSDF with voxel block grid (parallel spatial hashing)
  • Incremental 3D Gaussian Splatting. See here and MonoGS for a description of its backend [18].
Camera Settings
@@
  • Calibrate it and configure WEBCAM.yaml accordingly.
@@
Comparison pySLAM vs ORB-SLAM3
For a comparative evaluation, online trajectory estimated by pySLAM vs final trajectories estimated by ORB-SLAM3, see this nice notebook. Note that pySLAM is able to save both online and final pose estimates. On the other end, ORB-SLAM3 pose estimates are saved at the end of the full dataset playback. For further details about online/final trajectories and trajectory saving, see the section Trajectory saving.
Credits

### v3

@@ 4.4 Saving and reloading
When you run the script main_slam.py (main_map_dense_reconstruction.py):
  • You can save the current map state by pressing the button Save on the GUI. This saves the current map along with front-end, and backend configurations into the default folder results/slam_state (results/slam_state_dense_reconstruction).
  • To change the default saving path, open config.yaml and update target folder_path in the section:
[…]
    commandchars=
[…]
@@
[…]
    commandchars=
[…]
@@
[…]
    commandchars=
[…]
Note that pressing the Save button saves the current map, front-end, and backend configurations. Reloading a saved map replaces the current system configurations to ensure descriptor compatibility.
@@
Estimated trajectories can be saved in three formats: TUM (The Open Mapping format), KITTI (KITTI Odometry format), and EuRoC (EuRoC MAV format). pySLAM saves two types of trajectory estimates:
  • Online: In online trajectories, each pose estimate depends only on past poses. A pose estimate is saved at the end of each front-end iteration for the current frame.
  • Final: In final trajectories, each pose estimate depends on both past and future poses. A pose estimate is refined multiple times by LBA windows that include it, as well as by PGO and GBA during loop closures.
@@
[…]
commandchars=
[…]
Optimization engines
Currently, pySLAM supports both g2o and gtsam for graph optimization, with g2o set as the default engine. You can enable gtsam by setting to True the following parameters in config_parameters.py:
[…]
commandchars=
[…]
Additionally, the gtsam_factors package provides custom Python bindings for features not available in the original gtsam framework. See here for further details.
4.5 SLAM GUI
@@
  • Save: Save the map into the file map.json. You can visualize it back by using the script main_map_viewer.py (as explained above).
@@
  • Draw Ground Truth: If a ground truth dataset (e.g., KITTI, TUM, EUROC, or REPLICA) is loaded, you can visualize it by pressing this button. The ground truth trajectory will be displayed in 3D and will be progressively aligned with the estimated trajectory, updating approximately every 10-30 frames. As more frames are processed, the alignment between the ground truth and estimated trajectory becomes more accurate. After about 20 frames, if the button is pressed, a window will appear showing the Cartesian alignment errors along the main axes (i.e., $e_{x},e_{y},e_{z}$ and the history of the total $RMSE$ between the ground truth and the aligned estimated trajectories.
4.6 Monitor the logs for tracking, local mapping, and loop closing simultaneously
The logs generated by the modules local_mapping.py, loop_closing.py, loop_detecting_process.py, global_bundle_adjustments.py, and volumetric integrator_<X>.py are collected in the files local_mapping.log, loop_closing.log, loop_detecting.log, gba.log, and volumetric_integrator.log, which are all stored in the folder logs.
For debugging, you can monitor one of the them in parallel by running the following command in a separate shell:
@@ 4.6 Monitor the logs for tracking, local mapping, and loop closing simultaneously
Press CTRL+A followed by CTRL+Q to exit from tmux environment.
4.7 Evaluating SLAM
Run a SLAM evaluation
The main_slam_evaluation.py script enables automated SLAM evaluation by executing main_slam.py across a collection of datasets and configuration presets. The main input to the script is an evaluation configuration file (e.g., evaluation/configs/evaluation.json) that specifies which datasets and presets to be used. For convenience, sample configurations for the datasets TUM, EUROC and KITTI datasets are already provided in the evaluation/configs/ directory.
For each evaluation run, results are stored in a dedicated subfolder within the results directory, containing all the computed metrics. These metrics are then processed and compared. The final output is a report, available in PDF, LaTeX, and HTML formats, that includes comparison tables summarizing the Absolute Trajectory Error (ATE), the maximum deviation from the ground truth trajectory and other metrics.
Evaluation results are presented in the following subsections. For convenience, they are also available at this link. The evaluation uses commit f7c9a13 with the following presets: baseline, root_sift, and superpoint, which correspond to local feature extractors ORB2, root_SIFT, and SuperPoint, respectively.
TUM Evaluation
Table 1: TUM: Table RMSE
Dataset baseline root_sift superpoint
rgbd_dataset_freiburg1_desk 0.06634 0.05289 0.04815
rgbd_dataset_freiburg1_desk2 0.04645 0.06255 0.03868
rgbd_dataset_freiburg1_room 0.05885 0.07968 0.07199
rgbd_dataset_freiburg1_xyz 0.01196 0.01758 0.01726
rgbd_dataset_freiburg3_long_office_household 0.00921 0.00982 0.00952
rgbd_dataset_freiburg3_nostructure_texture_far 0.08397 0.11342 0.12118
rgbd_dataset_freiburg3_nostructure_texture_near_withloop 0.02245 0.05022 0.05037
Average 0.04275 0.05517 0.05102
Std Dev 0.0362 0.04531 0.04503
Best (Average) Preset baseline
Best (Average) Metric 0.04275
Table 2: TUM: Table Max
Dataset baseline root_sift superpoint
rgbd_dataset_freiburg1_desk 0.326 0.19361 0.22686
rgbd_dataset_freiburg1_desk2 0.44543 0.28865 0.38157
rgbd_dataset_freiburg1_room 0.14959 0.2838 0.17632
rgbd_dataset_freiburg1_xyz 0.06059 0.0677 0.0635
rgbd_dataset_freiburg3_long_office_household 0.03609 0.03543 0.04005
rgbd_dataset_freiburg3_nostructure_texture_far 0.29727 0.4361 0.70826
rgbd_dataset_freiburg3_nostructure_texture_near_withloop 0.09011 0.13774 0.11602
Average 0.20073 0.20615 0.24465
Std Dev 0.18661 0.18182 0.26334
Best (Average) Preset baseline
Best (Average) Metric 0.20073
Table 3: TUM: Table Percent Lost
Dataset baseline root_sift superpoint
rgbd_dataset_freiburg1_desk 1.364 1.502 1.674
rgbd_dataset_freiburg1_desk2 0.548 1.55 0.612
rgbd_dataset_freiburg1_room 0.074 0.148 0.088
rgbd_dataset_freiburg1_xyz 0.05 0.076 0.078
rgbd_dataset_freiburg3_long_office_household 0.0 0.0 0.0
rgbd_dataset_freiburg3_nostructure_texture_far 0.622 0.622 0.666
rgbd_dataset_freiburg3_nostructure_texture_near_withloop 0.024 0.036 0.036
Average 0.38314 0.562 0.45057
Std Dev 0.50438 0.82947 0.59847
Best (Average) Preset baseline
Best (Average) Metric 0.38314
EUROC Evaluation
Table 4: EUROC: Table RMSE
Dataset baseline root_sift superpoint
V101 0.08819 0.08816 0.08834
V102 0.06852 0.06609 0.06694
V201 0.07062 0.06409 0.09735
V202 0.62539 0.56925 0.53736
MH01 0.03775 0.03781 0.04172
MH02 0.04217 0.0456 0.04177
MH03 0.0512 0.04905 0.04882
MH04 0.0702 0.05907 0.0646
MH05 0.08177 0.06607 0.07131
Average 0.1262 0.11613 0.11758
Std Dev 0.19831 0.17517 0.15435
Best (Average) Preset root_sift
Best (Average) Metric 0.11613
Table 5: EUROC: Table Max
Dataset baseline root_sift superpoint
V101 0.16335 0.163 0.16208
V102 0.14545 0.12379 0.13136
V201 0.15864 0.18921 0.22163
V202 2.06607 2.52628 1.97847
MH01 0.09141 0.10259 0.21303
MH02 0.11922 0.1318 0.11196
MH03 0.19995 0.17533 0.14082
MH04 0.32021 0.27938 0.26722
MH05 0.24828 0.17474 0.18991
Average 0.39029 0.42957 0.37961
Std Dev 0.67128 0.75543 0.6093
Best (Average) Preset superpoint
Best (Average) Metric 0.37961
Table 6: EUROC: Table Percent Lost
Dataset baseline root_sift superpoint
V101 0.0 0.0 0.0
V102 0.0 0.0 0.0
V201 0.04 0.04 0.04
V202 0.146 0.138 0.112
MH01 0.05 0.074 0.08
MH02 0.018 0.026 0.012
MH03 0.046 0.03 0.032
MH04 0.0 0.01 0.0
MH05 0.07 0.016 0.034
Average 0.04111 0.03711 0.03444
Std Dev 0.05351 0.04829 0.04435
Best (Average) Preset superpoint
Best (Average) Metric 0.03444
KITTI Evaluation
Table 7: KITTI: Table RMSE
Dataset baseline root_sift superpoint
00 3.947 3.24849 3.93432
01 21.60014 21.86468 23.37292
02 6.1194 5.6707 10.14455
03 5.96451 5.92601 5.88895
04 1.74556 1.76121 1.65506
05 2.13915 2.11047 2.52263
06 2.74011 2.73981 2.6784
07 1.12358 1.12823 1.09496
08 3.99078 4.13702 4.14543
09 4.26068 4.35988 4.06244
10 2.22097 2.25127 2.27694
Average 5.07744 5.01798 5.61605
Std Dev 5.47265 5.63609 6.77
Best (Average) Preset root_sift
Best (Average) Metric 5.01798
Table 8: KITTI: Table Max
Dataset baseline root_sift superpoint
00 6.82983 5.8663 7.04067
01 33.32903 33.85867 36.62253
02 13.97623 9.47974 21.05188
03 10.16642 10.16978 10.06014
04 3.09135 3.0576 2.87271
05 3.81484 3.77687 4.9547
06 6.53758 5.14239 5.11673
07 1.84418 1.82355 1.8655
08 10.809 11.23635 10.75959
09 8.53888 8.70237 8.31166
10 4.13901 4.37849 4.27905
Average 9.37058 8.86292 10.26683
Std Dev 8.62548 8.62145 12.23529
Best (Average) Preset root_sift
Best (Average) Metric 8.86292
Table 9: KITTI: Table Percent Lost
Dataset baseline root_sift superpoint
00 0.0 0.038 0.012
01 0.0 0.0 0.0
02 0.012 0.004 0.004
03 0.0 0.0 0.0
04 0.0 0.0 0.0
05 0.0 0.0 0.028
06 0.108 0.018 0.0
07 0.0 0.0 0.0
08 0.0 0.0 0.0
09 0.0 0.0 0.0
10 0.0 0.0 0.0
Average 0.01091 0.00545 0.004
Std Dev 0.07244 0.02373 0.0196
Best (Average) Preset superpoint
Best (Average) Metric 0.004
pySLAM performances and comparative evaluations
For a comparative evaluation of the “online” trajectory estimated by pySLAM versus the “final” trajectory estimated by ORB-SLAM3, check out this nice notebook. For more details about “online” and “final” trajectories, refer to Section 4.4.
Note: Unlike ORB-SLAM3, which only saves the final pose estimates (recorded after the entire dataset has been processed), pySLAM saves both online and final pose estimates. For details on how to save trajectories in pySLAM, refer to this section. When you click the Draw Ground Truth button in the GUI (see here), you can visualize the Absolute Trajectory Error (ATE or RMSE) history and evaluate both online and final errors up to the current time.
5 Supported components and models
5.1 Supported local features
Figure 9: Timeline of some of the most famous local features for image matching, place recognition and SLAM.
Fig. 9 shows a timeline with some of the most famous local features for image matching, place recognition and SLAM. At present time, pySLAM supports the following feature detectors:
@@ 5.1 Supported local features
For more information, refer to pyslam/local_features/feature_types.py file. Some of the local features consist of a joint detector-descriptor. You can start playing with the supported local features by taking a look at test/cv/test_feature_manager.py and main_feature_matching.py.
In both the scripts main_vo.py and main_slam.py, you can create your preferred detector-descritor configuration and feed it to the function feature_tracker_factory(). Some ready-to-use configurations are already available in the file pyslam/local_features/feature_tracker.configs.py
The function feature_tracker_factory() can be found in the file pyslam/local_features/feature_tracker.py. Take a look at the file pyslam/local_features/feature_manager.py for further details.
@@
5.2 Supported matchers
See the file pyslam/local_features/feature_matcher.py for further details.
5.3 Supported global descriptors and local descriptor aggregation methods
@@
Different loop closing methods are available. These combines the above aggregation methods and global descriptors. See the file pyslam/loop_closing/loop_detector_configs.py for further details.
5.4 Supported depth prediction models
@@
5.5 Supported volumetric mapping methods
@@
  • TSDF with voxel block grid (parallel spatial hashing) [15]
  • Incremental 3D Gaussian Splatting. See here and MonoGS for a description of its backend [37, 21].
Supported semantic segmentation methods
  • DeepLabv3 [13]: from torchvision, pre-trained on COCO/VOC.
  • Segformer [70]: from transformers, pre-trained on Cityscapes or ADE20k.
  • CLIP [28]: from f3rm package for open-vocabulary support.
5.6 Configuration
Main configuration file
Refer to this section for how to update the main configuration file config.yaml and affect the configuration parameters in config_parameters.py.
Datasets
The following datasets are supported:
Dataset type in config.yaml
KITTI odometry data set (grayscale, 22 GB) type: KITTI_DATASET
TUM dataset type: TUM_DATASET
ICL-NUIM dataset type: ICL_NUIM_DATASET
EUROC dataset type: EUROC_DATASET
REPLICA dataset type: REPLICA_DATASET
TARTANAIR dataset type: TARTANAIR_DATASET
ScanNet dataset type: SCANNET_DATASET
ROS1 bags type: ROS1BAG_DATASET
ROS2 bags type: ROS2BAG_DATASET
Video file type: VIDEO_DATASET
Folder of images type: FOLDER_DATASET
Use the download scripts available in the folder scripts to download some of the following datasets.
KITTI Datasets
pySLAM code expects the following structure in the specified KITTI path folder (specified in the section KITTI_DATASET of the file config.yaml). :
[…]
commandchars=
[…]
  1. 1.
    Download the dataset (grayscale images) from http://www.cvlibs.net/datasets/kitti/eval_odometry.php and prepare the KITTI folder as specified above
  2. 2.
    Select the corresponding calibration settings file (section KITTI_DATASET: settings: in the file config.yaml)
TUM Datasets
pySLAM code expects a file associations.txt in each TUM dataset folder (specified in the section TUM_DATASET: of the file config.yaml).
  1. 1.
    Download a sequence from vision.in.tum.de/data/datasets/rgbd-dataset/download and uncompress it.
  2. 2.
    Associate RGB images and depth images using the python script associate.py. You can generate your associations.txt file by executing:
    \$ python associate.py PATH_TO_SEQUENCE/rgb.txt PATH_TO_SEQUENCE/depth.txt > associations.txt      # pay attention to the order!
  3. 3.
    Select the corresponding calibration settings file (section TUM_DATASET: settings: in the file config.yaml).
ICL-NUIM Datasets
Follow the same instructions provided for the TUM datasets.
EuRoC Datasets
  1. 1.
    Download a sequence (ASL format) from http://projects.asl.ethz.ch/datasets/doku.php?id=kmavvisualinertialdatasets (check this direct link)
  2. 2.
    Use the script io/generate_euroc_groundtruths_as_tum.sh to generate the TUM-like groundtruth files path + '/' + name + '/mav0/state_groundtruth_estimate0/data.tum' that are required by the EurocGroundTruth class.
  3. 3.
    Select the corresponding calibration settings file (section EUROC_DATASET: settings: in the file config.yaml).
Replica Datasets
  1. 1.
    You can download the zip file containing all the sequences by running:
    \$ wget https://cvg-data.inf.ethz.ch/nice-slam/data/Replica.zip
  2. 2.
    Then, uncompress it and deploy the files as you wish.
  3. 3.
    Select the corresponding calibration settings file (section REPLICA_DATASET: settings: in the file config.yaml).
Tartanair Datasets
  1. 1.
    You can download the datasets from https://theairlab.org/tartanair-dataset/
  2. 2.
    Then, uncompress them and deploy the files as you wish.
  3. 3.
    Select the corresponding calibration settings file (section TARTANAIR_DATASET: settings: in the file config.yaml).
ScanNet Datasets
  1. 1.
    You can download the datasets following instructions in http://www.scan-net.org. You will need to request the dataset from the authors.
  2. 2.
    There are two versions you can download:
    • A subset of pre-processed data termed as tasks/scannet_frames_2k: this version is smaller, and more generally available for training neural networks. However, it only includes one frame out of each 100, which makes it unusable for SLAM. The labels are processed by mapping them from the original Scannet label annotations to NYU40.
    • The raw data: this version is the one used for SLAM. You can download the whole dataset (TBs of data) or specific scenes. A common approach for evaluation of semantic mapping is to use the scannetv2_val.txt scenes. For downloading and processing the data, you can use the following repository as the original Scannet repository is tested under Python 2.7 and does’t support batch downloading of scenes.
  3. 3.
    Once you have the color, depth, pose, and (optional for semantic mapping) label folders, you should place them following {path_to_scannet}/scans/{scene_name}/[color, depth, pose, label]. Then, configure the base_path and name in the file config.yaml.
  4. 4.
    Select the corresponding calibration settings file (section SCANNET_DATASET: settings: in the file config.yaml). NOTE: the RGB images are rescaled to match the depth image. The current intrinsic parametes in the existing calibration file reflect that.
ROS1 bags
  1. 1.
    Source the main ROS1 setup.bash after you have sourced the pyslam python environment.
  2. 2.
    Set the paths and ROS1BAG_DATASET: ros_parameters in the file config.yaml.
  3. 3.
    Select/prepare the correspoding calibration settings file (section ROS1BAG_DATASET: settings: in the file config.yaml). See the available yaml files in the folder Settings as an example.
ROS2 bags
  1. 1.
    Source the main ROS2 setup.bash after you have sourced the pyslam python environment.
  2. 2.
    Set the paths and ROS2BAG_DATASET: ros_parameters in the file config.yaml.
  3. 3.
    Select/prepare the correspoding calibration settings file (section ROS2BAG_DATASET: settings: in the file config.yaml). See the available yaml files in the folder Settings as an example.
Video and Folder Datasets
You can use the VIDEO_DATASET and FOLDER_DATASET types to read generic video files and image folders (specifying a glob pattern), respectively. A companion ground truth file can be set in the simple format type: Refer to the class SimpleGroundTruth in io/ground_truth.py and check the script io/convert_groundtruth_to_simple.py.
5.7 Camera Settings
@@
  • Calibrate it and configure settings/WEBCAM.yaml accordingly.
@@
6 Credits